I don't know where to begin with this problem. can someone help me out?
Multiply. (11+3√3)(√11-3√3)

Answer by stanbon(75887)

Multiply. (11+3√3)(√11-3√3)
-----
Use FOIL:
F: 11*sqrt(11)
O: 11*-3sqrt(3) = -33sqrt(3)
I: 3sqrt(3)*sqrt(11) = 3sqrt(33)
L: 3sqrt(3)*-3sqrt(3) = -9*3 = -27
----
Combine those results to get:
-26+11Sqrt(11)-33sqrt(3)+3sqrt(33)
========================================
Comment:
I have a sneaking suspicion that you problem
is (sqrt(11)+3sqrt(3))(sqrt(11)-3sqrt(3))
---
If that is the case your answer is simply:
11-27 = -16
====================
Cheers,
Stan H.